OpenStudy (anonymous):

What is the standard form of y-8=2(x+3)?

OpenStudy (solomonzelman):

y-8=2x+6

OpenStudy (solomonzelman):

can you write this in a standard form?

OpenStudy (skullpatrol):

You need to begin with what is meant by "standard form"?

OpenStudy (skullpatrol):

@Titaniumxo in math you first must understand what is your "objective."

OpenStudy (anonymous):

When I have to change the form of the equation is where I mess up. I have y = 2x+14 but I need to have it to Ax+By=C. So, -2x + y = 14 2x - y = -14?

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

Thank you. :)

OpenStudy (skullpatrol):

Thanks for trying to understand :)